d/dt * t^2 =? a) 2t b) 4t^3 answer with explaination

d/dt * t^2 =? 
a) 2t         b) 4t^3
 answer with explaination

Grade:11

2 Answers

Shivansh Saumya
7 years ago
the formula for diff. such questions inovlving exponents and single variable is 
d(xn)/dx = n.xn-1
so, 
d(t2)/dt = 2.t2-1
 = 2.t1 = 2t
Anika Passi
5 years ago
We know that(d/dx)(xxn)=nnxn-1  {formula for differentiation)
So d(tt2)/dt= 2t2-1=2t1=2t
Thus A is correct option.
Hope it helps